OpenStudy (anonymous):

How do you reduce the order of a differential equation that has independent and dependent variables and no trivial solution? In my case: x''+4tx'+x^2=tan(t)t^2

OpenStudy (anonymous):

it is even analytically solvable?

OpenStudy (anonymous):

That's exactly the problem. I've been told that we need to replace x with two variables (y1 and y2). We then have 3 variables so I'm not really sure how it helps. We do however get a first order system. (y1)=x (y2)=x' (y3)=x''=(y2)' =>(y2)'+4t(y2)+(y1)^2=t^2tan(t)

OpenStudy (anonymous):

That would technically be first order but we haven't really done anything with the equation.

OpenStudy (anonymous):

If you can simplify it down to three first-order linear ODEs it can be solved using matrices and linear algebra.

OpenStudy (anonymous):

It becomes a eigenvalue problem.